QUESTION 8 To test whether the enactment of seatbelt legislation had any association with injuries to vehicle users, a random ****A Bit confused on this one- the original answer i put for the p value was 0.0000 but was incorrect- any input would help!*****
Answer #1

The p-value is 0.0000 (try putting it as this format 0.0000)

However, if it is not picking up, then you can try putting in more accurate value as 0.000000391

p=0.0000003912730495

Test Statistics
The Chi-Squared statistic is computed as follows:
\chi^2 = \sum_{i=1}^n \frac{(O_{ij}-E_{ij})^2}{E_{ij}}=32.5992


(4)P-value
The corresponding p-value for the test is p=Pr(χ2​>32.5992)=0


к M Chi-Square Independence test Observed Values RESULTS Total Post L 1320 l Prior L 5430 552 224 38 6750 622 Rows 4 Cols 2 d

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!